# Model Compression in Practice: Lessons Learned from Practitioners Creating On-device Machine Learning Experiences

### v1

@@ 5.6.2. Compounding Changes Degrade Accuracy
A common challenge with maintaining accuracy is that model compression techniques can compound in unintended ways. In some scenarios, small optimizations throughout a model build on one another, so by the time a data input reaches the end of a network, its prediction is totally off. In ML this phenomenon is similar to the concept of exploding/vanishing gradients (Bengio et al., 1994; Pascanu et al., 2013). The effect of this compounding error is not obvious to spot in beginning [E9].
@@
For quantization, practitioners advise a “gut check” to ensure the quantized weights and activations match the ranges of the original model, or else accuracy will be badly degraded [E6].
@@ 5.6.3. Robust, End-to-end Data Evaluation
Many teams do extensive model evaluation, error, and failure analysis [T2, E12, P11, P6, P7, E10, T4]. “We cannot assume compression doesn’t change model behavior, so we look at confusion matrices and instances where model gets it wrong” [E12]. In one project that P11, P6, and P7 worked on together where small mistakes could lead to a poor user experience, they built “unit tests” for different curated testing datasets to monitor their accuracy over time. E6 also emphasized using unit tests of data to observe how single instances move through the model to both understand how much computational change occurs after compression and make the necessary adjustments to the model code. “Data so influential to hitting accuracy targets” [E3].
@@ 5.6.4. Model Compression Artifacts
It is tempting to think of accuracy degradation as the only ML artifact; however, there are more subtle, even sinister implications depending on what a model is used for. For example, where some subpopulation of data is underrepresented in a dataset, e.g., at the tail of a distribution, it could be that ML compression techniques removes this tail, amplifying existing biases. Examples of this have been observed empirically in the few publications investigating robustness and evaluation of compressed models (Hooker et al., 2020, 2019; Ogueji et al., 2022; Liebenwein et al., 2021; Gale et al., 2019).
@@
One specific compression artifact example story was told from multiple participants. In the case of an object detection model that needed to run at a high frame rate, during development teams noticed one day that the bounding boxes were jittering in their demo. This finding was surprising, since the metrics from their most recent model iteration were reporting no changes, but “there were some weird side-effects” [T5]. It was not until someone debugging the problem realized that they had applied quantization throughout the neural network, including the final prediction layer. This coarsening of the data at the output produced correct bounding boxes, but resulted in a poor user experience. Another participant had seen this before, and remarked if the output of a model is “fine grain, don’t quantize” [T1].
Generalizing from this example, one learning that teams have found is that when a model’s prediction needs to be continuous, smooth, or user-perceivable at a high frame rate, it is best not to compress the final output layer (Deng et al., 2020). Participants told us that other modeling tasks can to be harder to optimize than classification [E5]. “Regression models have been an absolute nightmare to compress” [E7]. Anytime the output needs to be a continuous or floating point value, then compression techniques like quantization can produce artifacts [E12].
@@ 5.6.5. Demoing User Experiences
“Get [the experience] in hands of people as fast as possible.” — T4
@@
6. Design Opportunities for Efficient Machine Learning Tools and Interfaces
Given the unique challenges to efficient on-device ML, where can human-centered ML researchers, practitioners, and designers start to engage? As efficient machine learning techniques are driven forward by advances in hardware engineering and ML research, there remains a major barrier in practically applying these techniques for real-world ML features and user experiences. Tools influence what is possible: “there is a gap between what is possible with machine learning and what [tools] are being used,” [E4]. Currently, the tooling for efficient on-device ML is underdeveloped, “homegrown and ad-hoc,” [E9]. Moreover, current tools focus on individual algorithms (subsection 2.6) instead of the holistic process: “this a newer area, many tools are specific to a project and tend to be prototypes to prove feasibility,” [E9]. As we have demonstrated in this paper, there is considerable design planning, experimentation, and strategy that experts use to make on-device ML a reality. Proper tools could help guide and educate practitioners to develop these skills:
@@ 6.1. Developing Intuition for Compression
Although modern libraries for machine learning make it easier than ever develop models, that does not mean a practitioner know how to best train, evaluate, and deploy models. Machine learning is an inherently iterative and and empirical practice (Patel et al., 2008), and developing intuition for how models learn and behave is a major competitive advantage over blindly tweaking hyper-parameters. Interactive playgrounds that provide a safe environment where practitioners can quickly build and test their ideas could be a great onboarding experience for learning about efficient ML and model compression. Existing examples include TensorFlow Playground (Smilkov et al., 2016) for small neural networks, CNN Explainer (Wang et al., 2021c) for convolutional neural networks, and GANLab (Kahng et al., 2018) for generative additive models. An efficient ML playground could allow practitioners to learn about and build intuition about different compression techniques [E12], model architectures [E7], and their effect on hardware combinations [P9]. For example, comparing two compression techniques and inspecting their impact on metrics such as power, latency, and model size.
6.2. Comparing Models Across Compression Schemes
A common scenario in efficient model development is considering the trade-off between accuracy and performance, as discussed in Section 5.6. While conventional ML development requires model comparison between architectures and hyper-parameters, the additional optimization metrics that practitioners must navigate add additional complexity in efficient machine learning [T5]. Both accuracy and performance encompass numerous metrics to balance between. Better tooling to support model comparison could help practitioners compare a feasibility, or baseline model, against different compressed models to select the best model possible.
@@ 6.3. Finding Model Bottlenecks for Targeted Compression
In Section 5.5.5 and Section 5.6.4, practitioners discuss analyzing a model layer-by-layer. Targeted compression is the practice of selectively optimizing certain components of a model, for example, larger layers in a neural network. Practitioners described the operations that go into targeted compression. Currently the process is tedious, manual, and time-consuming, where “you must go layer by layer, operation by operation,” [P11, P6, P7]. There exists a space of interfaces that could help practitioners look inside their models, by mapping metrics to specific layers of a network [T5], finding the performance bottlenecks in a network [E5], comparing the input and output of these bottleneck layers [E9], and selectively optimizing them.
Targeted compression tooling may also ease the difficulty of thinking on the hardware level. For example, once an on-device model that is written in Python is compiled onto specialized hardware, many of the operations look far more low-level than Python code. To complicate matters, low-level hardware operations may not have a one-to-one mapping back to the original Python code, due to optimizations in the compilation process. Better tools could help practitioners perform analytics on their models to find bottlenecks while helping navigate between different layers of abstraction.
@@ 6.4. Evaluating Multi-model ML Systems
It is often assumed that model evaluation is done on a single model, isolated in a “lab environment.” In practice, this is not the case. Not only are models integrated into a larger app or codebase, many modern ML-powered experiences are the result of multiple models working together. This makes evaluation when models are arranged in chains, where the outputs of one model are the inputs of another [E10]. If you compress one model, how does that impact downstream models? Recalling the concept of compounding error discussed in Section 5.6, small errors introduced in earlier models can compound to produce bigger errors by the end of a model chain [E10]. E3 emphasized this challenge, saying that the future of ML driven user experiences will be accomplished through multi-model systems. When multiple models are running simultaneously, each can be evaluated individually, but also must be considered as a whole, which makes it “super hard to reason about,” [E3]. Future tools that generalize and can evaluate multi-model machine learning systems will have a big impact in helping practitioners build, debug, and make sense of large data-driven applications.
6.5. Easier Testing on Hardware Environments
One recommendation repeated again and again by practitioners in this study was to measure model performance on device. The performance of one model will like differ across different hardware [T4]. However, testing on real hardware can be a major barrier. Due the variety of mobile devices and different versions of hardware, it can be hard to build for multiple hardware implementations simultaneously. The average ML developer (outside of a hardware company) may not have access to all versions of all devices their users might have. When it comes to future tooling opportunities, E6 says it best, “the tools need to expose and take care of the cases for different hardware to maximize the hardware efficiency.”
@@ 6.6. Automating (Some) Compression Experiments
While many of the practitioners we interviewed discussed creating efficient models as an human-in-the-loop iteration process, there are opportunities to automate applying different compression schemes and present results to developers. An apt analogy is what AutoML is to hyperparameter searching: instead of sequentially trying different compression schemes, perform a grid search and parallelize many different tests simultaneously. P3 is excited by tools that could automate certain parts of a model, perform some analysis to figure out which techniques work and which do not, and finally present a summary of experiments make suggestions for the best fit compression schemes. However, we do not expect the process of creating efficient models to be completely automated. Throughout our study, experts made it clear that successful model compression is an iterative development process, but leveraging the strengths of automation where appropriate suggests rich mixed-initiative approaches to creating efficient models.
@@ 7. Limitations & Validity
Optimizing and creating efficient models to run on mobile devices is still relatively new, and best practices are evolving alongside the rapid pace of ML research. Thus we expect not all the practices mentioned in this paper will stand the test of time. As discussed in Section 4, studying experts from a single organization naturally limits the perspective of this paper, therefore we also expect certain details of our findings may not generalize. Nonetheless, we believe that the high-level strategies for on-device ML shared in this work—including budget design, strategies for investing in model optimization, and proper compression evaluation—will generalize to be useful for many different kinds of devices, domains, and ML user experiences. We are eager to watch how this interdisciplinary area of research progress in the future and reflect back on the practices outlined in this work to see which have remained, been removed, or reinvented.
@@ 8. Conclusion
In this research we illustrated some of the pragmatic design considerations that go into efficient on-device machine learning. While this paper is far less technical than related work from the ML literature, we feel the interdisciplinary bridge to UX and product design aspects are critical to bringing on-device ML into more approachable and popular practice. Through the results of this study, we are able to spotlight the holistic, end-to-end considerations that world-experts in model compression have developed from bridging ML research into real-life deployments of on-device ML experiences.

### v2

@@ 5.6.2. Compounding Changes Degrade Accuracy
A common challenge with maintaining accuracy is that model compression techniques can compound in unintended ways. In some scenarios, small optimizations throughout a model build on one another, so by the time a data input reaches the end of a network, its prediction is totally off. In ML, this phenomenon is similar to the concept of exploding/vanishing gradients (Bengio et al., 1994; Pascanu et al., 2013). The effect of this compounding error is not obvious to spot in beginning [E9]. For instance:
@@
For quantization, practitioners advise a “gut check” to ensure the quantized weights and activations match the ranges of the original model, or else accuracy will degrade [E6].
@@ 5.6.3. Robust, End-to-end Data Evaluation
Many teams do extensive model evaluation, error, and failure analysis [T2, E12, P11, P6, P7, E10, T4]. “We cannot assume compression doesn’t change model behavior, so we look at confusion matrices and instances where the model gets it wrong” [E12]. In one project that P11, P6, and P7 worked on where small mistakes could lead to a poor user experience, they built “unit tests” for different curated testing sets to monitor accuracy over time. E6 also emphasized using data unit tests to observe how single instances move through a model to both understand how much computational change occurs after compression and make the necessary adjustments to the model code. “Data so influential to hitting accuracy targets” [E3].
@@ 5.6.4. Model Compression Artifacts
It is tempting to think of accuracy degradation as the only ML artifact; however, there are more subtle, even sinister implications depending on what a model is used for. For example, where some subpopulation of data is underrepresented in a dataset, e.g., at the tail of a distribution, it could be that ML compression techniques remove this tail, amplifying existing biases. Examples of this have been observed empirically in the few publications investigating robustness and evaluation of compressed models (Hooker et al., 2020, 2019; Ogueji et al., 2022; Liebenwein et al., 2021; Gale et al., 2019).
@@
One specific compression artifact example story was told from multiple participants. In the case of an object detection model that needed to run at a high frame rate, during development teams noticed one day that the bounding boxes were jittering in their demo. This finding was surprising, since the metrics from their most recent model iteration were reporting no changes, but “there were some weird side-effects” [T5]. It was not until someone debugging the problem realized that they had applied quantization throughout the neural network, including the final prediction layer. This coarsening of the data at the output produced correct bounding boxes, but resulted in a poor user experience. Another participant had seen this before, and remarked if the output of a model is “‘fine grain,’ don’t quantize” [T1].
Generalizing from this example, one learning that teams have found is that when a model’s prediction needs to be continuous, smooth, or is user-perceivable at a high frame rate, it is best not to compress the final output layer (Deng et al., 2020). Participants told us that other modeling tasks can be harder to optimize than classification [E5]. “Regression models have been an absolute nightmare to compress” [E7]. Anytime the output needs to be a continuous or floating point value, compression techniques can produce artifacts [E12].
@@ 5.6.5. Demoing User Experiences
“Get [the user experience] in the hands of people as fast as possible.” — T4
@@
6. Design Opportunities for Efficient ML Tools and Interfaces
Given the unique challenges to efficient on-device ML, where can human-centered ML researchers, practitioners, and designers start to engage? As efficient machine learning techniques are driven forward by advances in hardware engineering and ML research, there remains a major barrier in practically applying these techniques for real-world features and user experiences.
Tools influence what is possible: “there is a gap between what is possible with machine learning and what [tools] are being used,” [E4]. Currently, the tooling for efficient on-device ML is underdeveloped, “homegrown and ad-hoc,” [E9]. Moreover, current tools focus on individual algorithms (Section 2.3.3) instead of the holistic process: “this is a newer area, many tools are specific to a project and tend to be prototypes to prove feasibility,” [E9]. As we have demonstrated in this paper, there is considerable design planning, experimentation, and strategy that experts use to make on-device ML a reality. Proper tools could help educate practitioners to develop these skills:
@@ 6.1. Developing Intuition for Compression
Although modern libraries for machine learning make it easier than ever develop models, that does not mean practitioners know how to best train, evaluate, and deploy models. Machine learning is an inherently iterative and and empirical practice (Patel et al., 2008; Amershi et al., 2019a), and developing intuition for how models learn and behave is a major competitive advantage over blindly tweaking hyper-parameters. Interactive playgrounds that provide a safe environment where practitioners can quickly build and test their ideas could be a great onboarding experience for learning about efficient ML and model compression. There is already precedent for this within machine learning, where interactive and educational tools help learners develop intuition around how certain models train and make predictions. Examples include include TensorFlow Playground (Smilkov et al., 2016) for small neural networks, CNN Explainer (Wang et al., 2021c) for convolutional neural networks, GANLab (Kahng et al., 2018) for generative adversarial models, and Diffusion Explainer (Lee et al., 2023) for stable diffusion models. These interactives typically run fully in-browser, enabling learners to access and experiment with ML techniques without installing software to needing access to extra compute. Moreover, they have been incorporated into ML curricula to help people gain complementary, hands-on experience for specifics model categories.
Imagine a playground for efficient ML, where practitioners could learn and build intuition about different compression techniques [E12], model architectures [E7], and their effect on hardware [P9]. Perhaps this playground could be web-based, where a small model is loaded in the browser running live inference. Users could select different compression techniques, with varying degrees of strength, and see the impact on the model, its metrics, and its predictions. This could help practitioners define realistic budgets (Section 5.4), test different combinations of architectures and compression techniques (Section 5.5.1), and inspect the difference between compute-bound or memory-bound models (Section 5.5.4).
6.2. Comparing Across Compression Schemes
A common scenario in efficient model development is considering the trade-off between accuracy and performance (Section 5.6.1). While conventional ML development requires model comparison between architectures and hyper-parameters, the optimization metrics that practitioners must also navigate add additional complexity when moving models on-device [T5]. Consider the scenario of having a well-performing model that does not hit a size budget. What do you do next? You can try every compression technique possible, but how do you ultimately select the best model that balances between the desired trade-offs? In a similar process discussed in Section 5.6.3, one practitioner mentioned that they “like to see an overview of a model’s robustness before and after compression,” [P8].
Better tooling to support model comparison could help practitioners compare a feasibility or baseline model (Section 5.4.1) against different compressed models to select the best model possible. There is rich opportunity for future work to investigate what to visualize and how. For example, flexible interfaces should handle visualizing the similarities and differences between models, their internals, their outputs, and applied compression techniques, where each technique not only has a suite of hyperparameters but can be also be combined with one another.
This line of work could also draw from existing work in experiment tracking. P10 described a project where the models they were developing were small, such that they could generate thousands of candidate models. This participant always compared candidates against an uncompressed baseline model, and maintained documentation tracking their experimental history, with a few notes per model, so that other stakeholders could make an informed model selection. However, this required diligent effort by the developer—experimental histories tend to exist across multiple documents and can result in long tables of values where patterns can be hard to discover. Future opportunities for tooling point to model tracking systems that can generate interactive reports to help practitioners observe their model development history, while helping them select the best model (from potentially thousands of models) that passes their accuracy threshold and maximizes performance.
@@ 6.3. Finding Model Bottlenecks for Targeted Compression
In Section 5.5.4 and Section 5.6.4, practitioners discuss analyzing a model layer-by-layer. Targeted compression is the practice of selectively optimizing specific components of a model, for example, individual layers in a neural network. Currently, practitioners described the process as tedious, manual, and time-consuming, where “you must go layer by layer, operation by operation,” [P11, P6, P7]. There exists a space of interfaces that could help practitioners look inside their models, by mapping metrics to specific layers of a network [T5], finding the performance bottlenecks in a network [E5], comparing the input and output of these bottleneck layers [E9], and selectively optimizing them.
Targeted compression tooling may also ease the difficulty of thinking at the hardware level. E6 gave an example: once an on-device model written in Python is compiled onto specialized hardware, the operations are expressed by what the hardware can support, and is likely not as readable as documented Python code. To complicate matters, low-level hardware operations may not have a one-to-one mapping back to the original Python code, due to optimizations in the compilation process. Better tools could help practitioners perform analytics on their models to find bottlenecks and easily traverse between different layers of abstraction (e.g., model code and hardware operations).
@@ 6.4. Evaluating Multi-model ML Systems
It is often assumed that model evaluation is done on a single model, isolated in a “lab environment.” In practice, this is not the case. Not only are models integrated into larger apps or codebases, many modern ML-powered experiences are the result of multiple models working together. For example, models could be arranged in chains, where the outputs of one model are the inputs of another [E10]. If you compress one model, how does that impact downstream models? Recalling the concept of compounding error discussed in Section 5.6.2, small errors introduced in earlier models can compound to produce bigger errors by the end of a model chain [E10]. E3 emphasized this challenge, saying that the future of ML driven user experiences will be accomplished through multi-model systems. When multiple models are running simultaneously, each can be evaluated individually, but also must be considered as a whole, which makes it “super hard to reason about,” [E3]. Future tools that generalize and can evaluate multi-model machine learning systems will have a big impact in helping practitioners build, debug, and make sense of large data-driven applications.
6.5. Simplified Hardware Testing
One recommendation repeated by practitioners in this study was to measure model performance on device (Section 5.5.2). T4 emphasized that the performance of one model will like differ across different hardware (Section 5.5.3). However, testing on real hardware can be a major barrier. Due the variety of mobile devices and different versions of hardware, it can be hard to build for multiple hardware implementations simultaneously. The average ML developer (outside of a hardware company) may not have access to all versions of all devices their users might have. When it comes to future tooling opportunities, E6 says it best, “the tools need to expose and take care of the cases for different hardware to maximize the hardware efficiency.” While there is undoubtedly room for ML-hardware innovation here, for the purposes of this work assume this problem can be reduced to looping over a set of hardware and evaluating a model. With all these results, this problem could be cast as yet another comparison task, where tool designers need to enhance existing workflows to allow ML practitioners to track, organize, and see what hardware passed or failed certain criteria.
@@ 6.6. Automating (Some) Compression Experiments
While many of the practitioners we interviewed discussed creating efficient models as a human-in-the-loop iteration process, there are opportunities to automate applying different compression schemes and present results to developers. An apt analogy is what AutoML is to hyperparameter searching: instead of sequentially trying different compression schemes, perform a grid search and parallelize many different tests simultaneously. P3 was particularly excited about compression automation. Imagine mixed-initiative tools that take in a model and user-specified budgets, runs through a suite of compression techniques to find all possible experiments that satisfy the budgets, and finally presents a summary of recommended compression recipes for a practitioner to choose from. Perhaps through tools like these we could even learn something about the behavior of compression itself. However, we do not expect the process of creating efficient models to be completely automated. Throughout our study, experts made it clear that successful model compression is an iterative development process (Section 5.2), but leveraging the strengths of automation where appropriate suggests rich opportunity for mixed-initiative approaches for creating efficient models.
@@ 7. Limitations & Validity
Optimizing and creating efficient models to run on mobile devices is still relatively new, and best practices are evolving alongside the rapid pace of ML research. Thus we expect not all the practices mentioned in this paper will stand the test of time. As discussed in Section 4, studying experts from a single organization naturally limits the perspective of this paper, therefore we also expect certain details of our findings may not generalize. Nonetheless, we believe that the high-level strategies for on-device ML shared in this work—including budget design, strategies for investing in model optimization, and proper compression evaluation—will generalize to be useful for many different kinds of devices, domains, and ML user experiences. We are eager to watch how this interdisciplinary area of research progresses in the future and reflect back on the practices outlined in this work to see which have remained, which have been removed, and which have been reinvented.
@@ 8. Conclusion
In this research we illustrated some of the pragmatic design considerations that go into efficient on-device machine learning. While this paper is far less technical than related work from the ML literature, we feel the interdisciplinary bridge to UX and product design are critical to bringing on-device ML into more approachable and popular practice. Through the results of this study, we are able to spotlight the holistic, end-to-end considerations that experts in model compression have developed from translating ML research into intelligent, on-device ML experiences.